OpenStudy (anonymous):

what is the arc length when theta= pi/3 and the radius is 5 cm? a. 10pi/3 cm b. 16pi/3 cm c. pi/3 cm d. 5pi/3 cm

OpenStudy (anonymous):

SOMEONE HELP MEEEE PLEASE D;

OpenStudy (anonymous):

s = r*theta

OpenStudy (anonymous):

s = 5*pi/3

OpenStudy (anonymous):

I dont think I did it right, what answer did you get?

OpenStudy (anonymous):

What answer did you get?

OpenStudy (anonymous):

I guess d, Im not sure

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

D is correct. since arc length = radius * angle in radians

OpenStudy (anonymous):

ok thank you. do you think you could help me with another?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

the function f(t)= 25 sin(pi/2t) +10 models the temperature of a periodic chemical reaction where t represent time in hours. What are the maximum and minimum temperatures of the reaction, and how long does the entire cycle take?

OpenStudy (anonymous):

is is it f(t) = 25*sin((pi/2)*t) + 10, or f(t) = 25*sin((pi/2*t)) + 10

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

well. We know the sin function has values between -1 and 1.

OpenStudy (anonymous):

And, since we have 25*sin(pi/2 * t) + 10, we know -15 <= f(t) <= 35

OpenStudy (anonymous):

Do you see how we got that?

OpenStudy (anonymous):

okay, so what do we do now? I understand how you got that but Im completely clueless with these types of problems ;(

OpenStudy (anonymous):

all this problem is asking for is the range and the periodicity of f(t). We just stated the range, and now we must state the periodicity. Our independent variable, t, is scaled by pi/2. What does this mean? Basically, our function has a period twice as long as sin(t) would. Check out this site and play around with the applets towards the bottom to get a better idea of how to handle these questions. http://www.intmath.com/trigonometric-graphs/3-graphs-sin-cos-phase-shift.php

OpenStudy (anonymous):

okay I will, so did you get an answer for this? I can give you the options if you want me to

OpenStudy (anonymous):

can you at least tell me the maximum and minimum?

OpenStudy (anonymous):

Sure, go ahead and list them. We've already stated the answer.

OpenStudy (anonymous):

-15 <= f(t) <= 35

OpenStudy (anonymous):

oh okay the answer was A! thank you. this is the same type of problem, can you please help? this is the last one The function f(t)= 3 cos(pi/6t) +5 represents the tide in the Blastic Sea. It has a maximum of 8 feet when time (t) is 0 and a minimum of 2 feet. The sea repeats this cycle every 12 hours. After nine hours, how high is the tide? a. 12 feet b. 5 feet c. 4.5 feet d. 2.5 feet

OpenStudy (anonymous):

it's not actually (pi/6t) its more like (pi/6) *t) like the other problem

OpenStudy (anonymous):

Ok, so the idea here is relate 12 hours to the period of the function.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Are the answers listed correct?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Its 5. Sorry for the delay. i must go now.

OpenStudy (anonymous):

okay thank you so much!